The odds of an event not occuring are given as 2 to 2. What is the probability of success?
(Enter exact fraction or a decimal answer to a minimum of 4 places.)
Answer #1

ODDS of an event not occuring is 2 : 2

(As odd is the no. of events happening / no. o events not happening)

P( Success) = 2/ (2+2) = 2/4 = 0.5